OpenStudy (anonymous):

what is the solution of |x-4|-10=1

OpenStudy (saifoo.khan):

|x-4| = 11

OpenStudy (saifoo.khan):

x-4 = +11 & x-4 = -11 solve.

OpenStudy (saifoo.khan):

x = 15 & x = -7

OpenStudy (anonymous):

\[|x-4|-10=1\] \[|x-4|=11\] because of the bars, |x-4| can either be positive of negative when it's positive, x-4=11, so x = 15 when it's negative x-4 =-11 so x = -7